Solve the system:
x-2y=1
2x+3y=23
I get it, but the answer is not coming out for me :( please help me :)

Answer by algebrahouse.com(1659)

x - 2y = 1 ------> x = 2y + 1 {added 2y to both sides in top equation}
2x + 3y = 23

2x + 3y = 23 {original bottom equation}
2(2y + 1) + 3y = 23 {substituted 2y + 1, in for x, into bottom equation}
4y + 2 + 3y = 23 {used distributive property}
7y + 2 = 23 {combined like terms}
7y = 21 {subtracted 2 from both sides}
y = 3 {divided both sides by 7}

x = 2y + 1 {re-arranged top equation}
x = 2(3) + 1 {substituted 3, in for y, into 2y + 1}
x = 6 + 1 {multiplied 2 by 3}
x = 7 {added 6 and 1}

x = 7 and y = 3

For more help from me, visit: www.algebrahouse.com